OpenStudy (anonymous):

pleasseee help!! Graph the function. y = -x + x2

OpenStudy (anonymous):

wait, is that \[y=-x + x^2 ?\]

OpenStudy (anonymous):

is it point up or down

OpenStudy (anonymous):

This is a difference of two squares problem

OpenStudy (anonymous):

for \[ax^2+bx+c \] the value of a determines whether it opens up or down. (if a is positive then it opens up...)

OpenStudy (anonymous):

is more of it on the first quadrant then the second?

OpenStudy (anonymous):

More? parabolas extend to infinity...

OpenStudy (anonymous):

like to the left or right the majoriy of the openness

OpenStudy (anonymous):

the majority of the graph will be in the first quadrant.

OpenStudy (anonymous):

\(\large y=-x+x^2=x(x-1) \) written in standard form, \(\large y=ax^2+bx+c=\color {red}1\cdot x^2-\color {red} 1 \cdot x+\color {red}0 \) easy to see the x intercepts are at x=0, x=1, the vertex is x=\(\large -\frac{b}{2a} \) where a=1, b=-1... this is a parabola opening upwards (a is positive)